# LayerSkip: Enabling Early Exit Inference and Self-Speculative Decoding

### v2

@@ 2.1 Exiting Earlier in LLMs
Similar analysis was done in Geva et al. (2022) on a GPT2 model Radford et al. (2019) as it developed predictors to estimate when prediction saturates to exit early. For the particular example we present in Figure 2, we find, on average, a token requires 23.45 layers out of the model’s 32 layers. Hence, even if we have a perfect predictor that has zero compute overhead, we can only save up to 26% of computation. Therefore, there is a need to make LLM models require fewer layers to predict each token, and spend less compute being hesitant or “changing its mind”. By default, deep learning models are not motivated to predict their final output early and instead spread their compute across all layers Voita et al. (2019, 2023). We see in Figure 2(b), that tokens we would consider easy or straightforward to predict, e.g., Token 02 that starts a for-loop, required all 32 layers to predict “for”. We would like our model to be less reliant on later layers and only use later layers for harder tokens. We would like our models to be more reliant on earlier layers than later layers. To do that, we propose skipping layers during training, which we refer to as layer dropout. However, we use higher dropout rates for later layers and lower dropout rates for earlier layers, to make the model less reliant on later layers.
@@ 3 Related Work
Elbayad et al. (2020) added a dedicated LM head for each decoder layer in an encoder-decoder translation model, and explored different early exit granularities (per token as well as per sequence) in both training and inference, and explored exiting early both statically and dynamically, using auxiliary layers that learned during training when to exit. CALM Schuster et al. (2022) built upon  Elbayad et al. (2020) and started with a model pretrained with early exit losses, and focused on finding optimal criteria to decide which layer to exit at during inference. SkipDecode Corro et al. (2023) fine-tuned a decoder-only model to skip more layers for later tokens within a sequence.  Din et al. (2023) started with pretrained models and finetuned auxiliary fully-connected layers to map the embeddings outputted by earlier layers to later layers. In our proposed solution, we do not introduce any additional modules or linear layers for early exit, and instead used a shared exit for all layers.
@@ 4.1 Training using Layer Dropout & Early Exit Loss
We denote the input tokens to a transformer model as $X$ and its output as $Y$, with an embedding layer that maps the token indices to token embeddings, $x_{0}$, and a transformer model with $L$ transformer layers, where transformer layer $l$ evolves embeddings outputted from its previous layer, $x_{l+1}=x_{l}+f_{l}(x_{l})$, and a final LM head that maps the embedding outputs of the last layer, $x_{L}$ to logits, $e_{L}=g(x_{L})$. We denote the binary cross entropy loss function that is usually used to train language models as $J_{\text{BCE}}(e_{L},Y)$.
@@ 5 Experiments
  • Continual Pretraining: start with a pretrained model and continue pretraining on 52B tokens from a corpus of diverse data containing natural language text and code. We experiment using pretrained Llama2 7B (32 layers), with $p_{\text{max}}=0.1$, $e_{scale}=0.2$, $C_{\text{rot},R=8}$, and Llama2 13B (40 layers), with $p_{\text{max}}=0.1$, $e_{scale}=0.1$, $C_{\text{rot},R=39}$.
@@ 6.2 Self-Speculative Decoding Results
We evaluate the self-speculative decoding algorithm introduced in §4.3 on the different trained models. We report quality metrics, EM (exact match) and ROUGE-2 Ganesan (2018) implemented in torchmetrics Nicki Skafte Detlefsen et al. (2022), token acceptance rate for the self speculation algorithm (how often verification accepts each of the draft tokens), throughput measured as tokens per token averaged over the sampled dataset, and finally a speed up compared to autoregressive decoding. For our early exit and our self-speculative decoding experiments, we specify the layer we exit at in the $E$ column in each table below. We compare with Draft & Verify Zhang et al. (2023) on common models and tasks evaluated in both papers. All experiments were performed with greedy decoding and generated a maximum of 512 tokens for each sample. Following  Zhang et al. (2023), speedup is calculated as the acceleration of average inference time per token compared to the “Autoregressive” baseline on the same setting. “Autoregressive” experiments use baseline models that were pretrained or finetuned without LayerSkip, while “Early Exit” and “Self Speculative” experiments use our models trained or finetuned with LayerSkip. Our implementation leverages HuggingFace Wolf et al. (2020).
Continual Pretraining In Table 6.2, we evaluate the continual pre-training of Llama2 7B and 13B with and without LayerSkip on various tasks: CNN/DM Nallapati et al. (2016), XSUM Narayan et al. (2018) abstractive summarization tasks, and HumanEval Chen et al. (2021) coding task. The experiments were performed on NVIDIA H100 GPUs. The number of speculations, i.e., the number of tokens generated in the draft stage in self-speculation, is specified in the $d$ column. We follow Zhang et al. (2023) and perform CNN/DM with 1-shot abstractive summarization, and XSUM with 0-shot. We obtain speedups between 1.34$\times$ and and 2.16$\times$ depending on the model or the task. In general, we observe higher speedups for the smaller Llama2 7B compared to the larger Llama2 13B model. When comparing with Draft & Verify Zhang et al. (2023) on the common model and tasks of both approaches, we are significantly faster on CNN/DM (1.81$\times$ vs. 1.5$\times$) and slightly slower on XSUM (1.34$\times$ vs. 1.48$\times$).
Pretraining from Scratch Our results are presented in Table 6.2. The experiments were performed on NVIDIA H100 GPUs. We found that an opposite tried to continual pretraining: the bigger model has a bigger speedup than the smaller model, reaching 2.16$\times$ speedup, which exceeds traditional speculative decoding.
@@ 7 Ablation Studies
Scaling with Pretraining Tokens In order to understand how the accuracy of last and middle layers change across time when pretraining from scratch, we ran 3 training experiments with different number of tokens on Llama 1.5B and show the results in Figure 11. Each experiment trained for 50,000 steps, per device batch size of 4, context window of 4096, but changed the number of GPUs to 32, 64, 128. We plotted the perplexity of a held out split of The Stack on the last layer (layer 24) and the middle layer (layer 12). As expected, perplexity on last layer decreases as we train on more tokens. However, surprisingly, we discover that perplexity on middle layer increases drastically by default in training, unless we apply early exit loss. Layer dropout reduces the increase as well. This could open the door to more research on the dynamics of transformers and the evolution of embeddings in earlier layers to understand why embeddings across layers are close to each other early on in training but diverge drastically as training progresses. This could also present a motivation for our training recipe that has minimal drop in last layer accuracy while significantly improves accuracy of earlier layers.
@@ Acknowledgements
We would like to thank Volker Seeker, Artem Korenev, and Ilia Kulikov for logistic support, FAIR’s clusters’ support team members, especially, Henry Estela, Hongsheng Song, Shubho Sengupta, and Nabib Ahmed, for their help in maintaing our clusters, Kamila Benzina for legal support, Fabian Gloeckle, Andrey Gromov, Francisco Massa, Daniel Haziza, Aaditya Singh, Karen Hambardzumyan, Nicola Cancedda, for discussions.

### v4

@@ 2.1 Exiting Earlier in LLMs
Similar analysis was done in Nostalgebraist (2020) and  Geva et al. (2022) on a GPT2 model Radford et al. (2019) as it developed predictors to estimate when prediction saturates to exit early. For the particular example we present in Figure 2, we find, on average, a token requires 23.45 layers out of the model’s 32 layers. Hence, even if we have a perfect predictor that has zero compute overhead, we can only save up to 26% of computation. Therefore, there is a need to make LLM models require fewer layers to predict each token, and spend less compute being hesitant or “changing its mind”. By default, deep learning models are not motivated to predict their final output early and instead spread their compute across all layers Voita et al. (2019, 2023). We see in Figure 2(b), that tokens we would consider easy or straightforward to predict, e.g., Token 02 that starts a for-loop, required all 32 layers to predict “for”. We would like our model to be less reliant on later layers and only use later layers for harder tokens. We would like our models to be more reliant on earlier layers than later layers. To do that, we propose skipping layers during training, which we refer to as layer dropout. However, we use higher dropout rates for later layers and lower dropout rates for earlier layers, to make the model less reliant on later layers.
@@ 3 Related Work
Elbayad et al. (2020) added a dedicated LM head for each decoder layer in an encoder-decoder translation model, and explored different early exit granularities (per token as well as per sequence) in both training and inference, and explored exiting early both statically and dynamically, using auxiliary layers that learned during training when to exit. CALM Schuster et al. (2022) built upon  Elbayad et al. (2020) and started with a model pretrained with early exit losses, and focused on finding optimal criteria to decide which layer to exit at during inference. SkipDecode Corro et al. (2023) fine-tuned a decoder-only model to skip more layers for later tokens within a sequence.  Din et al. (2023) started with pretrained models and finetuned auxiliary fully-connected layers to map the embeddings outputted by earlier layers to later layers.  Zeng et al. (2023) and  Raposo et al. (2024) added trainable routers to determine which layers to skip during inference during pretraining. In our proposed solution, we do not introduce any additional modules or linear layers for early exit, and instead used a shared exit for all layers.
@@ 4.1 Training using Layer Dropout & Early Exit Loss
We denote the input tokens to a transformer model as $X$ and its output as $Y$, with an embedding layer that maps the token indices to token embeddings, $x_{0}$, and a transformer model with $L$ transformer layers, where transformer layer $l$ evolves embeddings outputted from its previous layer, $x_{l+1}=x_{l}+f_{l}(x_{l})$, and a final LM head that maps the embedding outputs of the last layer, $x_{L}$ to logits, $e_{L}=g(x_{L})$. We denote the cross entropy loss function that is usually used to train language models as $J_{\text{CE}}(e_{L},Y)$.
@@ 5 Experiments
  • Continual Pretraining: start with a pretrained model and continue pretraining on 52B tokens from a corpus of diverse data containing natural language text and code. We experiment using pretrained Llama2 7B (32 layers), with $p_{\text{max}}=0.1$, $e_{scale}=0.2$, $C_{\text{rot},R=8}$, and Llama2 13B (40 layers), with $p_{\text{max}}=0.1$, $e_{scale}=0.1$, $C_{\text{rot},R=39}$. We later continually pretrained Llama3 8B on 419B tokens with $p_{\text{max}}=0.1$, $e_{scale}=0.1$, $C_{\text{rot},R=8}$, and and Llama3.2 1B on 839B tokens with $p_{\text{max}}=0.1$, $e_{scale}=0.1$, $C_{\text{rot},R=8}$.
@@ 6.1 Early Exit Inference Results
For Llama3 8B and Llama 3.2 1B, we notice a larger drop in last layer accuracy when continually pretraining with LayerSkip, despite training on more tokens. The perplexity of earlier layers of the baseline Llama3 models are 2 or 3 orders of magnitude higher than their Llama2 counterparts. Reducing such high perplexity of earlier layers while maintaining accuracy of last layer is more challenging and could be the cause of the drop in last layer accuracy. The increase in perplexity of earlier layers of Llama3 compared to Llama2 could be due to the increase of number of pretraining tokens (8T for Llama3 versus 2T for Llama2), as the ablation in Section 7 confirms. This could be a motivation to consider our LayerSkip recipe in pretraining future LLMs from scratch.
{NiceTabular}
[…]
Last Layer Middle Layer Last Layer Middle Layer
(Layer 32) (Layer 16) (Layer 16) (Layer 8)
Eval Perplexity
[…]
Common Sense Reasoning Tasks
(Multiple Choice Questions / True False Questions)
BoolQ 83.3 84.0 62.3 79.5 64.2 63.5 62.2 57.7
PIQA 79.0 79.0 59.9 71.3 75.0 72.9 56.7 62.8
SIQA 45.5 45.0 35.7 42.0 42.0 42.3 33.8 37.2
HellaSwag 60.0 59.2 28.0 44.8 47.2 45.5 26.5 31.7
WinoGrande 72.8 71.7 55.4 65.8 61.3 60.2 50.0 51.9
ARC-e 78.7 79.4 33.9 62.7 67.1 66.8 28.3 41.7
ARC-c 51.0 48.2 27.2 31.9 33.0 32.3 23.9 20.2
OBQA 34.0 33.6 19.0 21.6 28.2 26.6 19.6 18.4
COPA 89 87 58 81 79 79 58 62
Reading Comprehension
(Multiple Choice Questions)
RACE Middle 62.3 61.1 25.9 51.5 51.6 52.5 22.7 32.1
RACE High 44.9 43.1 24.8 36.6 36.7 37.1 33.8 37.2
MMLU
(Multiple Choice Questions)
MMLU 66.5 60.5 27.6 38.1 31.2 24.8 25.8 25.9
Question Answering
(Open Ended Answers)
NaturalQuestions 30.3 27.5 0.00 4.63 11.9 9.06 0.00 1.08
TriviaQA 65.5 62.5 0.00 14.8 36.1 31.3 0.00 1.95
Mathematics
(Open Ended Answers)
GSM8K 54.2 45.0 0.00 2.27 5.84 3.79 0.00 2.05
MATH 17.3 12.3 0.00 0.72 1.70 1.78 0.00 0.04
Code Generation
(Open Ended Answers)
HumanEval 37.8 28.7 0.00 7.32 17.7 9.15 0.00 1.22
MBPP 49.0 40.0 0.00 13.6 25.8 13.2 0.00 0.00
Table 2: Evaluation of continual pretraining of Llama3 8B and Llama3.2 1B.
@@ 6.2 Self-Speculative Decoding Results
We evaluate the self-speculative decoding algorithm introduced in §4.3 on the different trained models. We report quality metrics, EM (exact match) and ROUGE-2 Ganesan (2018) implemented in torchmetrics Nicki Skafte Detlefsen et al. (2022), token acceptance rate for the self speculation algorithm (how often verification accepts each of the draft tokens), throughput measured as tokens per second averaged over the sampled dataset, and finally a speed up compared to autoregressive decoding. For our early exit and our self-speculative decoding experiments, we specify the layer we exit at in the $E$ column in each table below. We compare with Draft & Verify Zhang et al. (2023) on common models and tasks evaluated in both papers. All experiments were performed with greedy decoding and generated a maximum of 512 tokens for each sample. Following  Zhang et al. (2023), speedup is calculated as the acceleration of average inference time per token compared to the “Autoregressive” baseline on the same setting. “Autoregressive” experiments use baseline models that were pretrained or finetuned without LayerSkip, while “Early Exit” and “Self Speculative” experiments use our models trained or finetuned with LayerSkip. Our implementation leverages HuggingFace Wolf et al. (2020).
Continual Pretraining In Table 6.2, we evaluate the continual pre-training of Llama2 7B and 13B with and without LayerSkip on various tasks: CNN/DM Nallapati et al. (2016), XSUM Narayan et al. (2018) abstractive summarization tasks, and HumanEval Chen et al. (2021) coding task. The experiments were performed on NVIDIA H100 GPUs. The number of speculations, i.e., the number of tokens generated in the draft stage in self-speculation, is specified in the $d$ column. We follow Zhang et al. (2023) and perform CNN/DM with 1-shot abstractive summarization, and XSUM with 0-shot. We obtain speedups between 1.34$\times$ and 2.16$\times$ depending on the model or the task. In general, we observe higher speedups for the smaller Llama2 7B compared to the larger Llama2 13B model. When comparing with Draft & Verify Zhang et al. (2023) on the common model and tasks of both approaches, we are significantly faster on CNN/DM (1.81$\times$ vs. 1.5$\times$) and slightly slower on XSUM (1.34$\times$ vs. 1.48$\times$).
Pretraining from Scratch Our results are presented in Table 6.2. The experiments were performed on NVIDIA H100 GPUs. We found that an opposite trend to continual pretraining: the bigger model has a bigger speedup than the smaller model, reaching 2.16$\times$ speedup, which exceeds traditional speculative decoding.
@@ 7 Ablation Studies
Scaling with Pretraining Tokens In order to understand how the accuracy of last and middle layers change across time when pretraining from scratch, we ran 3 training experiments with different number of tokens on Llama 1.5B and show the results in Figure 11. Each experiment trained for 50,000 steps, per device batch size of 4, context window of 4096, but changed the number of GPUs to 32, 64, 128. We plotted the perplexity of a held out split of The Stack dataset on the last layer (layer 24) and the middle layer (layer 12). As expected, perplexity on last layer decreases as we train on more tokens. However, surprisingly, we discover that perplexity on middle layer increases drastically by default in training, unless we apply early exit loss. Layer dropout reduces the increase as well. This could open the door to more research on the dynamics of transformers and the evolution of embeddings in earlier layers to understand why embeddings across layers are close to each other early on in training but diverge drastically as training progresses. This could also present a motivation for our training recipe that has minimal drop in last layer accuracy while significantly improves accuracy of earlier layers.
@@ Acknowledgements
We would like to thank Volker Seeker, Artem Korenev, and Ilia Kulikov for logistic support, FAIR’s clusters’ support team members, especially, Henry Estela, Hongsheng Song, Shubho Sengupta, and Nabib Ahmed, for their help in maintaing our clusters, Helen Klein, Kamila Benzina, and Ty Toledano for legal support, Fabian Gloeckle, Andrey Gromov, Francisco Massa, Daniel Haziza, Aaditya Singh, Karen Hambardzumyan, Nicola Cancedda, for discussions.